# Completely regular growth solutions to linear differential equations with exponential polynomials coefficients

### v1

@@ 4 Proofs
$\sum_{l=1}^{K}e^{c_{l}t+O(\log|t|)}\cdot\sum_{m=0}^{n}\hat{R}_{l,j}(t)(1+o(1))\frac{g^{(j)}}{g}=o\left(e^{ar}\right),\quad r=|t|,\enspace K\in\mathbb{N},\enspace\in\mathbb{C}.$ (5.5)
@@
for $t\in S_{m}(\varepsilon)$, every $C>0$, $r\geq r_{0}$. Thus, the remanent proof is the same with the sections in 3.2 and 3.6. The part relevant to denote the angular domains is more complicated, which is omitted here. Besides some areas near the critical rays, and parabolic strips, it is always established the fundamental coefficient differential equations are in the form of
@@
By the similar proof of theorem 3.2 with Lemma 2.4, it follows
@@
We set the Class of finite order transcendental solutions of (3.6) in Theorem 3.6, and (5.2) in Theorem 5.1 as $\mathcal{EP}^{0}$, $\mathcal{EP}^{1}$ individually. On this point of view, if we take the solutions of equation (5.2) as coefficients of a linear differential again, then we will get the same conclusion the finite order transcendental solutions are of c.r.g. in $\mathcal{EP}^{2}$ and the orders are rational. As if we do this progress recursively, even if there may exist new finite order transcendental functions, but every asymptotic form in a narrow angular domain with an open angle $\varepsilon$ is just the same form as
@@
From this property, it is natural to explore, what is the maximal probability of the class of functions of exponential polynomial type and if the functional class will be closed. By comparison between the function of exponential polynomial type (5.8), and the definition of c.r.g. in an angular domain in the form of
$\log|f|=h(\theta)r^{\rho}+o(r^{\rho}),$
we find that there may exist commonality. When $\rho$ is a rational number, their difference is just the small quantity of $r^{\rho}$, referring to the equality
$h(\theta)=A(\theta)\cos(\rho\theta+\varphi),$
for $z=re^{i\theta}$ in a narrow angular domain $S$, where the density of zeros in $S$ is zero, and $A,\varphi\in\mathbb{R}$ in the book of Levin[9, p.115]. This property is also very natural for function of exponential polynomial type in $S$ on the rays where zeros not mainly distributed, from the conclusion that the zeros here are just $O(\log r)$ resulting from Steinmetz [12, Cor. 1].
On the other, it leaves some space for us to reflect on:
@@
    How large is the gap between the two classes of functions, of rational order with completely regular growth, and the one of exponential polynomial type?

### v2

@@ 5.1 A nested concept – exponential polynomial type
$\sum_{l=1}^{K}e^{c_{l}t+O(\log|t|)}\cdot\sum_{m=0}^{n}\hat{R}_{l,j}(t)(1+o(1))\frac{g^{(j)}}{g}=o\left(e^{ar}\right),\quad r=|t|,\kern 5.0ptK\in\mathbb{N},\kern 5.0pt\in\mathbb{C}.$ (5.5)
@@
for $t\in S_{m}(\varepsilon)$, every $C>0$, $r\geq r_{0}$. Thus, the remanent proof is the same with the sections in 3.2 and 3.9. The part relevant to denote the angular domains is more complicated, which is omitted here. Besides some areas near the critical rays, and parabolic strips, it is always established the fundamental coefficient differential equations are in the form of
@@
By the similar proof of Theorem 3.2 with Lemma 2.4 (when Lemma 2.4 lose effect for infinite many zeros of $f$ in each sector, we take the same discussion with (4.10) into account, which is abbreviated in this part), it follows
@@
We set the Class of finite order transcendental solutions of (3.8) in Theorem 3.9, and (5.2) in Theorem 5.1 as $\mathcal{EP}^{0}$, $\mathcal{EP}^{1}$ individually. On this point of view, if we take the solutions of equation (5.2) as coefficients of a linear differential again, then we will get the same conclusion the finite order transcendental solutions are of c.r.g. in a class defined as $\mathcal{EP}^{2}$ and the orders are rational. As if we do this progress recursively, even if there may exist new finite order transcendental functions, but every asymptotic form in a narrow sector with an open angle $\varepsilon$ is just the same form as
@@
This corollary reveals a fundamental closure property: the class of functions of exponential polynomial type, defined inductively as $\bigcup_{k=0}^{\infty}\mathcal{EP}^{k}$, constitutes a vast and coherent hierarchy in which the property of completely regular growth is hereditary. Starting from the classical exponential polynomials ($\mathcal{EP}^{0}$), each step $\mathcal{EP}^{k}\to\mathcal{EP}^{k+1}$ corresponds to taking all finite-order transcendental solutions of linear differential equations whose coefficients lie in $\mathcal{EP}^{k}$. Our main theorems guarantee that this process never leaves the world of completely regular growth.
While the inductive definition suggests a potentially expanding hierarchy, the strict inclusion $\mathcal{EP}^{k}\subsetneq\mathcal{EP}^{k+1}$ is not established rigorously here. Perturbative constructions, such as those alluded to in later sections, might generate new functions from $\mathcal{EP}^{0}$ to $\mathcal{EP}^{1}$ for an example, hinting that the classes are indeed proper extensions. However, a full verification would require explicit examples and lies beyond the immediate scope. This hierarchical perspective naturally raises deeper questions about the scope and limits of the class of exponential polynomial type, especially in relation to the broader theory of completely regular growth.
5.2 An Operator-Theoretic Perspective and Parameter Space Analysis
The classical question of existence and regularity of finite-order solutions to linear differential equations can be reframed fruitfully within the framework of operator theory on function spaces. This perspective not only clarifies the underlying mechanisms of growth inheritance but also opens pathways to construct new classes of solutions and to analyze the structure of the parameter space in perturbed equations.
The asymptotic theory of linear differential equations naturally leads to integral formulations. Consider a homogeneous linear differential equation
$f^{(n)}+P_{1}f^{(n-1)}+\cdots+P_{n-1}f^{\prime}+P_{n}f=0.$
By introducing suitable approximating functions $Q_{i}$ and defining the non-homogeneous term
$F(f)=(Q_{1}-P_{1})f^{(n-1)}+\cdots+(Q_{n}-P_{n})f,$
the equation can be rewritten as
$Lf:=f^{(n)}+Q_{1}f^{(n-1)}+\cdots+Q_{n}f=F(f).$
Using the method of variation of constants, the general solution takes the form
$f(z)=\sum_{i=1}^{n}c_{i}F_{i}(z)+\sum_{i=1}^{n}F_{i}(z)\int_{g_{i}}^{z}\frac{\Delta_{i}(t)}{\Delta(t)}F(f(t))dt,$
where $\{F_{i}\}$ are linearly independent solutions of $Lf=0$, $\Delta$ is the Wronskian, and $\Delta_{i}$ are corresponding cofactors [36, p. 128]. This representation is a Volterra integral equation of the second kind:
$f(z)=f_{\text{hom}}(z)+\int_{0}^{z}K(z,t)f(t)dt,$ (5.10)
where $f_{\text{hom}}$ is a particular solution of the homogeneous equation $Lf=0$ and $K$ is a kernel derived from the coefficients. The associated affine integral operator
$(Tf)(z)=\tilde{f}(z)+\int_{0}^{z}K(z,t)f(t)dt$ (5.11)
becomes the central object of study.
To control growth, we work in weighted Banach spaces of entire functions. For a continuous, non-increasing weight $v:[0,\infty)\to[0,\infty)$ with $\lim_{r\to\infty}r^{m}v(r)=0$ for all $m\in\mathbb{N}$, define
$\mathcal{H}_{v}^{\infty}(\mathbb{C}):=\{f\in\mathcal{H}(\mathbb{C}):\|f\|_{v}:=\sup_{z\in\mathbb{C}}v(|z|)|f(z)|<\infty\}.$
Of particular importance are the exponential weights $v(r)=e^{-ar^{\sigma}}$ ($a>0$, $\sigma>0$), which give spaces
$\mathcal{B}_{\sigma,a}=\left\{f\in\mathcal{H}(\mathbb{C}):\|f\|_{\sigma,a}=\sup_{z\in\mathbb{C}}|f(z)|e^{-a|z|^{\sigma}}<\infty\right\}.$
These spaces are closely related to generalized weighted Bergman spaces of entire functions; see, e.g., [3, 5, 10, 11, 25, 26]. An entire function $f$ belongs to $\mathcal{B}_{\sigma,a}$ if its growth order $\rho(f)$ satisfies $\rho<\sigma$, or $\rho=\sigma$ with type $\tau<a$.
Within this framework, the inheritance of completely regular growth (c.r.g.) for equations with exponential polynomial coefficients can be understood as a consequence of the compression properties of the associated Volterra operator on these weighted spaces. The structure of the coefficients ensures the integral kernel preserves the geometric constraints needed for the operator to act as a contraction, forcing the iterative scheme to converge to a well-behaved (c.r.g.) solution.
The integral operator formulation offers a potential pathway for constructing solutions and analyzing their growth. Consider a perturbed equation of the form
$L(f)=\varepsilon F(f),$ (5.12)
where $L$ and $F$ are differential polynomials with polynomial coefficients and exponential polynomial coefficients seperately, and $\deg L>\deg F$. Under this framework, the operator $T$ in (5.11) can be formally decomposed as $Tf=f_{\text{hom}}+\varepsilon T^{\prime}f$. A natural line of inquiry is to investigate the conditions under which $T^{\prime}$ acts boundedly on a suitably chosen weighted Bergman space $\mathcal{B}_{\sigma,a}$. If such boundedness can be established, then for sufficiently small $|\varepsilon|$, the operator $\varepsilon T^{\prime}$ might become a contraction, leading to a unique fixed point of $T$ via the Banach fixed-point theorem. This fixed point would then correspond to a solution of the differential equation with controlled growth properties.
This perspective, while promising, immediately raises several non-trivial questions and highlights the delicate interplay between the operator’s properties, the choice of function space, and the perturbation parameter:
@@ 5.2 An Operator-Theoretic Perspective and Parameter Space Analysis
    Feasibility of the boundedness condition. Under what precise conditions on the coefficients of $L$ and $F$ is the associated operator $T^{\prime}$ bounded on a space $\mathcal{B}_{\sigma,a}$? The answer is not automatic and depends crucially on the structure of the integral kernel derived from the differential equation.